(3.3.19) Use the following cell phone airport data speeds​ (Mbps) from a particular network. Find the percentile corresponding to the data speed 0.8 Mbps
0.1
0.1
0.3
0.3
0.5
0.6
0.6
0.6
0.7
0.7
0.8
0.8
0.8
0.9
0.9
1.3
1.4
1.5
1.6
1.8
2.1
2.2
2.2
2.7
2.8
2.9
3.4
3.5
4.7
4.8
5.8
6.4
6.7
8.4
9.8
10.8
11.1
11.8
11.9
12.4
12.5
12.6
12.9
13.3
13.4
13.8
13.8
14.8
15.1
27.4
Percentile of 0.8=
1
Expert's answer
2020-12-06T20:27:09-0500

Sorted data:



percentile = (number of values less than x) / (total number of values) * 100% = 10/50*100% = 20%